What is the price today (in dollars and cents) of a 15-year zero coupon bond if the required rate of return is 8.99%. The bond face value is $1000. $ Place your answer in dollars and cents. You should set your calculator for at least four decimal places of accuracy.
Answer #1

Price = FV / ( 1 + r)n

Price = 1,000 / ( 1 + 0.0899)15

Price = 1,000 / 3.637473

Price = $274.9161

You can also use a financial calculator:

FV 1000

N 15

I/Y 8.99

CPT PV